A metal has a density of 10.22 g/cm^3, atomic weight of 95.94 g/mole and atomic radius of 0.136 nm. Is it BCC or FCC?
1
Expert's answer
2020-11-20T08:06:56-0500

Density ("\rho") = 10.22"g\/cm^3"

Atomic weight (A) = 95.94g/mol

Atomic radius (R) = 0.136nm = 1.36 × "10^{-8}"cm



In order to determine whether metal has a FCC or a BCC crystal structure we need to compute the density for each of the crystal structures.


FCC ---->


"\rho = \dfrac{nA}{(2\sqrt{2}R)^3N_A}"


"\rho = \dfrac{4×95.94}{(2\sqrt{2}×1.36×10^{-8})^3× 6.02× 10^{23}}"


"= 11.19g\/cm^3"



BCC ----->


"\rho = \dfrac{nA}{(\frac{4}{\sqrt{3}}R)^3N_A}"


"\rho = \dfrac{2×95.94}{(\frac{4}{\sqrt{3}}×1.36×10^{-8})^3× 6.02×10^{23}}"


"= 10.22g\/cm^3"



The density derived from the BBC formula is the value we're provided with in the problem. Therefore, the metal has a BCC crystal structure, and not a FCC structure.